# Observational prospects of self-interacting scalar superradiance with next-generation gravitational-wave detectors

### v1

@@ 1 Introduction
While the superradiant instability only relies on gravity, interactions among the ultralight particles and possible couplings to the Standard Model or other sectors can drastically alter the evolution of the cloud of bound states [44, 66, 67, 68]. While these can be model-specific, both the QCD axion and the ultralight scalars predicted from higher-dimensional theories generically exhibit some level of self-interactions. Effects of self-interactions on scalar cloud evolution have been considered since early on [18, 69, 33], but energy exchange processes between different bound states were largely neglected. Earlier, the belief was that strong attractive self-interactions lead to the collapse of the entire cloud back into the BH, a phenomenon dubbed a “bosenova” [18]. However, as intuited first with a toy model by [70] and later analyzed in detail for the full dynamical evolution of a BH-scalar system in [44], self-interactions lead to energy exchange between different bound states, the BH, and infinity. Equilibrium is generally reached at bound-state occupations smaller than those required for the violent “bosenova”. It is now well-established through both analytic analysis and numerical simulations [44, 71] that relatively strong self-interactions of the bosonic particles can lead to steady-states with orders of magnitude smaller occupation numbers than the expectation from purely gravitational growth. This prevents BH spin-down, keeps the cloud below the bosenova threshold, and suppresses GW emission.
@@ 2.1 The superradiant instability
There are also corrections to the real parts of the frequencies due to attractive gravitational self-interactions of the cloud, which make it more bound. These are important for the calculation of frequency drifts and were computed in [29, 44]. For simplicity, we use 211 to denote the $n=2,l=1,m=1$ level; similar for 322. For the 211 and 322 levels, which we will see are the most significant ones in section 2.3, these corrections are
@@ 2.2 Self-interacting scalars
The QCD axion is a well-known example of such a particle, generated at low energies as a result of a Peccei-Quinn (PQ) symmetry broken at a high energy scale $f_{a}$. At energies below the QCD scale, the axion acquires a symmetric potential. Because the low-energy axion potential is generated by QCD, its mass $m_{b}$ and the PQ-breaking scale $f_{a}$ are related by $m_{b}c^{2}\simeq 6\times 10^{-12}\,{\rm{eV}}(10^{18}\,{\rm{GeV}}/f_{a})$, resulting in quartic coupling uniquely fixed by the particle mass,
@@ 2.3 Superradiance of self-interacting scalar fields
Energy and angular momentum conservation between the two levels and the BH lead to a set of coupled first-order differential equations for the (normalized) occupation number of the levels, $\varepsilon_{nlm}\equiv N_{nlm}\hbar c/(GM^{2})$, the dimensionless BH spin, $\chi$, and the mass of the BH, $M$:
@@
We denote the rates in the form $\gamma_{\rm{initial\,state}}^{\rm{final\,state}}$. In Equations (9) and (10), the first line corresponds to gravitational processes, namely superradiance and GW emission, while the second line corresponds to non-relativistic energy exchange processes between the different levels (see Figure 1).
@@ 2.4 Regimes of scalar superradiance
Here we outline the behaviour of self-interacting superradiant clouds in the $(m_{b},f)$ plane. Four main regimes are described in Ref. [44], in the order of increasing self-interaction strength: the gravitational regime, the moderate self-interaction regime, the strong self-interaction regime, and the no spin-down regime. The first two, the gravitational and moderate self-interaction regimes, are demonstrated in this paper to be of particular observational interest. We subsequently focus on the distinct dynamics and signatures of only these two regimes. A detailed analysis of all four regimes can be found in Ref. [44].
@@
it takes roughly $\log(J/(m\hbar))\sim 180$ e-folds of superradiant growth to backreact on the spacetime. As a result, in the steady-state configuration, the BH will have been spun down by an $\mathcal{O}(1)$ fraction, and the cloud will have a high occupation number $\sim\mathcal{O}(J/\hbar)$. To leading order in $\alpha$, the final BH spin after saturation of the 211 level is $\chi\approx\chi_{i}-4\alpha/(1+4\alpha^{2})$, where $\chi_{i}$ is the initial spin of the BH. The energy density stored in a cloud around a stellar-mass BH can be as large as that of a white dwarf, $\sim 10^{9}$ kg/m3, so any pre-existing scalar abundance in the neighborhood of the BH does not appreciably change either the final occupation number or the required e-folding time.
The bosonic cloud has an oscillating quadrupole moment, which sources gravitational radiation. The emission occurs at an angular frequency of $\sim 2m_{b}c^{2}/\hbar$, which can heuristically be thought of as boson annihilations. Because of the non-relativistic nature of the hydrogenic-like clouds, GW power $P$ is a steep function of $\alpha$, $P\propto GN^{2}m_{b}^{4}\alpha^{4l+16}\hbar^{2}c^{3}$, where $N$ is the occupation number of the cloud. Nevertheless, the decay timescale can be short compared to astrophysical timescales for stellar-mass BHs, which manifests as cloud depletion over time.
The next level to grow will reach its maximum occupation after $\sim 180$ e-folds of its superradiant growth. As the BH is spun down to an even smaller spin, the previous level becomes unstable and gets re-absorbed. Because of the hierarchy of superradiant timescales, the previous level has been depleted appreciably due to GW emission when it falls back into the BH. The angular momentum of the black hole continues to decrease and the mass can temporarily increase as the now unstable level gets absorbed. Thus the evolution proceeds in the same spirit, until astrophysical timescales become relevant, such as the BH’s age or accretion rate. A representative time evolution of the occupation numbers and BH mass and spin is shown in the upper panels of Figure 2.
@@
The moderate self-interaction regime: The crossover from the gravitational regime is characterized by two conditions: 1) the age of the BH, $T_{\rm{BH}}$, is large enough to allow the process $211\times 211\to 322\times{\rm BH}$ to appreciably populate 322, and 2) the first level, 211, has not been significantly depleted by GW emission by this time. Combining these two conditions, the moderate self-interaction regime occurs for [44]
@@
A representative time evolution of the occupation numbers and BH mass and spin is shown in the lower panels of Figure 2. Here, the 211 level grows unimpeded to its maximum occupation number, $\varepsilon_{211}\approx 1$, and spins down the BH as in the purely gravitational evolution. Because of the stronger self-interactions, the 322 level can grow much faster than $\sim 180$ e-folds of its superradiant growth. As the 322 level grows to an appreciable size, re-pumping of the 211 level through $322\times 322\to 211\times\infty$ becomes significant and the cloud enters a regime where the ratio $\varepsilon_{211}/\varepsilon_{322}>1$ remains constant, forming a quasi-equilibrium. The two-level cloud loses energy to the BH and to infinity, primarily through quartic processes.
Since the 322 level grows through self-interactions to a very large occupation, its superradiant growth drives it to maximum occupation in just a few e-folds, thus reaching saturation much earlier than in the purely gravitational growth regime. As shown in the rightmost panels of Figure 2, for moderate self-interactions, 322 has grown maximally in $\sim 10^{6}$ yr (bottom row) compared to $\sim 10^{8}$ yr in the purely-gravitational growth regime (top row).
Once the quartic process $211\times 211\to 322\times{\rm BH}$ becomes faster than the superradiance of 211, the initial growth of 211 is suppressed by a simultaneous growth of 322. In this case, we have a much smaller maximum occupation of the 211 level, $\varepsilon_{211}^{\max}\ll 1$, and it occurs for [44]
@@
Breakdown of perturbation theory: In all of the aforementioned regimes, the perturbative treatment of self-interactions is justified in the sense that the hydrogenic wavefunctions are not significantly distorted, timescales for all rates are much longer than oscillation times, and the dynamics are restricted to a two-level system. For $\alpha\gtrsim 0.22$, the two-level approximation no longer holds and higher levels can grow. While perturbative evolution can, in principle, be hierarchical and closed in non-gravitational regimes for $\alpha\lesssim 0.3$ [44], no demonstration of a closed regime which is under good analytic control has been performed for $\alpha\gtrsim 0.3$. We do not include these high-$\alpha$ scenarios in our current work as we only have control of our approximations in the purely gravitational growth regime, where $f$ is close to the Planck scale. There are potential signatures in the regime of $\alpha\gtrsim 0.22$, as shown in the results of this study, so we stress the importance of studying the high-$\alpha$ dynamics and parameter space in the future.
@@ 3.2 Occupation numbers and gravitational-wave strain
The emission frequencies for the 211 and 322 annihilations are broadly similar for a given $m_{b}$, whereas the transition signal is at a much lower frequency. For annihilation signals, the GW frequency corresponds to twice the energy of the annihilated particles, whereas for transition signals, it is the difference between the energy levels. The energy of the 211 and 322 levels differs only to the second order in $\alpha$ [see Equation (2.1)], and thus for small $\alpha$ as considered here, the energy of the 211 and 322 levels differ by a small relative amount. This leads to similar annihilation frequencies but a much lower transitional energy difference than the total particle energy, leading to low-frequency transitional GWs.
@@ 4.2 Observable parameter space
Figure 7: Parameter space reachable by current and next-generation detectors (enclosed by the solid curves; see legend) for (left) 211 annihilation, (middle) 322 annihilation, and (right) $322\rightarrow 211$ transition signals from an ultralight scalar cloud around Cygnus X-1, assuming parameters as in Table 2 and an age of $1.0\times 10^{5}$ yr. The dashed curves divide the parameter spaces into three regimes: (top) the strong self-interaction regime, (middle) the moderate self-interaction regime, and (bottom) the gravitational regime. The reachable regions for Einstein Telescope (enclosed by black curve) and Cosmic Explorer (enclosed by red curve) largely overlap. The parameter space reachable by MAGIS-Space is limited to the brightest spot in the right panel in this example. In other cases it can probe a larger parameter space; see for, e.g., MOA-2011 in Figure 11.
@@ 5 Conclusions
We note that in this work we have only considered central measured values on some BH parameters and assumptions on others (like the initial spin and age) that have large uncertainties, so our constraints cannot be considered as true exclusions in the scalar particle parameter space. GW searches marginalized over the BH uncertainties, in tandem with more precise measurements of the BH properties, would be of great interest.
@@ C.1 Cygnus X-1 with an age of $10^{6}$ yr
Figure 10: (Similar to Figure 7) Parameter space reachable by current and next-generation detectors (enclosed by the solid curves; see legend) for signals from Cygnus X-1, assuming parameters as in Table 2 and an age of $1.0\times 10^{6}$ yr. The reachable regions for the 211 annihilation, 322 annihilation, and $322\rightarrow 211$ transition signals are shown in the left, middle, and right panels, respectively. The dashed curves divide the parameter spaces into three regimes: (top) the strong self-interaction regime, (middle) the moderate self-interaction regime, and (bottom) the gravitational regime. The reachable regions for Einstein Telescope (enclosed by black curve) and Cosmic Explorer (enclosed by red curve) largely overlap. The transition signal is not reachable by MAGIS-Space for this particular example.

### v2

@@ 1 Introduction
While the superradiant instability only relies on gravity, interactions among the ultralight particles and possible couplings to the Standard Model or other sectors can drastically alter the evolution of the cloud of bound states [66, 44, 67, 68, 69, 70, 71, 72, 73]. While these can be model-specific, both the QCD axion and the ultralight scalars predicted from higher-dimensional theories generically exhibit some level of self-interactions. Effects of self-interactions on scalar cloud evolution have been considered since early on [18, 74, 33], but energy exchange processes between different bound states were largely neglected. Earlier, the belief was that strong attractive self-interactions lead to the collapse of the entire cloud back into the BH, a phenomenon dubbed a “bosenova” [18]. However, as intuited first with a toy model by [66] and later analyzed in detail for the full dynamical evolution of a BH-scalar system in [44], self-interactions lead to energy exchange between different bound states, the BH, and infinity. Equilibrium is generally reached at bound-state occupations smaller than those required for the violent “bosenova”. It is now well-established through both analytic analysis and numerical simulations [44, 72] that relatively strong self-interactions of the bosonic particles can lead to steady-states with orders of magnitude smaller occupation numbers than the expectation from purely gravitational growth. This prevents BH spin-down, keeps the cloud below the bosenova threshold, and suppresses GW emission.
@@ 2.1 The superradiant instability
There are also corrections to the real parts of the frequencies due to attractive gravitational self-interactions of the cloud, which make it more bound. These can be computed by solving for the gravitational potential induced by the cloud through the Poisson equation, and then for the induced energy shifts to the energy eigenstates characterized by the appropriate $(n,l,m)$ numbers. Energy corrections from self-gravity depend on the occupation number of the cloud, and so are important for the calculation of frequency drifts. The computation was carried out in [29, 44]. For simplicity, we use 211 to denote the $n=2,l=1,m=1$ level; similar for 322. For the 211 and 322 levels, which we will see are the most significant ones in section 2.3, these corrections are
@@ 2.2 Self-interacting scalars
The QCD axion is a well-known example of such a particle, generated at low energies as a result of a Peccei-Quinn (PQ) symmetry broken at a high energy scale $f_{a}$. At energies below the QCD scale, the axion acquires a symmetric potential. Because the low-energy axion potential is generated by QCD, its mass $m_{b}$ and the PQ-breaking scale $f_{a}$ are related by $m_{b}c^{2}\simeq 6\times 10^{-12}\,{\rm{eV}}(10^{18}\,{\rm{GeV}}/f_{a})$, resulting in quartic coupling uniquely fixed by the particle mass [83],
@@ 2.3 Superradiance of self-interacting scalar fields
In perturbation theory, where the spectrum does not get modified by self-interactions, the dynamics of the system can be fully characterized by the evolution of the occupation numbers. This is described by a set of coupled first-order ordinary differential equations that follow directly from energy and angular momentum conservation between the different levels, the absorbing BH and infinity, and were derived in [44]. The equations for the (normalized) occupation number of the levels, $\varepsilon_{nlm}\equiv N_{nlm}\hbar c/(GM^{2})$, the dimensionless BH spin, $\chi$, and the mass of the BH, $M$ are:
@@
We denote the rates in the form $\gamma_{\rm{initial\,state}}^{\rm{final\,state}}$. In particular, $\gamma_{\rm BH}^{nlm}$ describes the standard superradiant rate for the $nlm$ level and $\gamma_{nlm\times nlm}^{\rm GW}$ describes annihilations to GWs from the level $nlm$. Furthermore, $\gamma_{nlm\times nlm}^{n^{\prime}l^{\prime}m^{\prime}\times\infty}$ describes the energy transfer of two quanta from the $nlm$ level to one quantum of the $n^{\prime}l^{\prime}m^{\prime}$ level and to one quantum of radiation to infinity. $\gamma_{nlm\times nlm}^{n^{\prime}l^{\prime}m^{\prime}\times{\rm BH}}$ describes a similar process where, instead, one quantum of energy falls back into the BH. As such, in Equations (9) and (10), the first line corresponds to gravitational processes, namely superradiance and GW emission, while the second line corresponds to non-relativistic energy exchange processes between the different levels (see Figure 1).
@@ 2.4 Regimes of scalar superradiance
Here we outline the behaviour of self-interacting superradiant clouds in the $(m_{b},f)$ plane. Four main regimes are described in Ref. [44], in the order of increasing self-interaction strength: the gravitational regime, the moderate self-interaction regime, the strong self-interaction regime, and the no spin-down regime. The first two, the gravitational and moderate self-interaction regimes, are demonstrated in this paper to be of particular observational interest. Example evolutions of systems in these regimes are shown in the upper and lower panels of Figure 2, respectively, and we here focus on the distinct dynamics and signatures of only these two regimes. A detailed analysis of all four regimes can be found in Ref. [44].
@@
it takes roughly $\log(J/(m\hbar))\sim 180$ e-folds of superradiant growth to backreact on the spacetime. As a result, in the steady-state configuration, the BH will have been spun down by an $\mathcal{O}(1)$ fraction, and the cloud will have a high occupation number $\sim\mathcal{O}(J/\hbar)$. This is seen in the upper panels of Figure 2 at a time of $\sim 10^{0}$ yr. To leading order in $\alpha$, the final BH spin after saturation of the 211 level is $\chi\approx\chi_{i}-4\alpha/(1+4\alpha^{2})$, where $\chi_{i}$ is the initial spin of the BH. The energy density stored in a cloud around a stellar-mass BH can be as large as that of a white dwarf, $\sim 10^{9}$ kg/m3, so any pre-existing scalar abundance in the neighborhood of the BH does not appreciably change either the final occupation number or the required e-folding time.
The bosonic cloud has an oscillating quadrupole moment, which sources gravitational radiation. The emission occurs at an angular frequency of $\sim 2m_{b}c^{2}/\hbar$, which can heuristically be thought of as boson annihilations. Because of the non-relativistic nature of the hydrogenic-like clouds, GW power $P$ is a steep function of $\alpha$, $P\propto GN^{2}m_{b}^{4}\alpha^{4l+16}\hbar^{2}c^{3}$, where $N$ is the occupation number of the cloud [33]. Nevertheless, the decay timescale can be short compared to astrophysical timescales for stellar-mass BHs, which manifests as cloud depletion over time.
The next level to grow will reach its maximum occupation after $\sim 180$ e-folds of its superradiant growth. As the BH is spun down to an even smaller spin, the previous level becomes unstable to decay and gets re-absorbed. This is seen in the upper panels of Figure 2 at a time of $\sim 10^{8}$ yr. Because of the hierarchy of superradiant timescales, the previous level has been depleted appreciably due to GW emission when it falls back into the BH. The angular momentum of the black hole continues to decrease and the mass can temporarily increase as the now unstable level gets absorbed. Thus the evolution proceeds in the same spirit, until astrophysical timescales become relevant, such as the BH’s age or accretion rate.
@@
The moderate self-interaction regime: The crossover from the gravitational regime is characterized by two conditions: 1) the age of the BH, $T_{\rm{BH}}$, is large enough to allow the process $211\times 211\to 322\times{\rm BH}$ to appreciably populate 322, and 2) the first level, 211, has not been significantly depleted by GW emission by this time. This computation has been carried out in Section IV.B.3 of [44]. Schematically, the first condition amounts to using Equation (10) and demanding that the self-interactions-induced process $\gamma_{211\times 211}^{322\times{\rm BH}}(\varepsilon_{211}^{\max})^{2}\gtrsim\log(GM^{2})/T_{\rm BH}$, namely the time it takes to build up 322 in the lifetime $T_{\rm BH}$ of the BH. The second condition similarly comes about from Equation (9) and is the requirement $\gamma_{211\times 211}^{\rm GW}\varepsilon_{211}\lesssim\gamma_{211\times 211}^{322\times{\rm BH}}(\varepsilon_{211}^{\max})^{2}/\log(GM^{2})$. Combining these two conditions, the moderate self-interaction regime occurs for
@@
A representative time evolution of the occupation numbers and BH mass and spin is shown in the lower panels of Figure 2. Here, the 211 level grows unimpeded to its maximum occupation number, $\varepsilon_{211}\approx 1$, and spins down the BH as in the purely gravitational evolution. Because of the stronger self-interactions, the 322 level can grow much faster than $\sim 180$ e-folds of its superradiant growth. As the 322 level grows to an appreciable size, re-pumping of the 211 level through $322\times 322\to 211\times\infty$ becomes significant and the cloud enters a regime where the ratio $\varepsilon_{211}/\varepsilon_{322}>1$ remains constant, forming a quasi-equilibrium. This quasi-equilibrium can be seen in the lower panels of Figure 2 for times between $\sim 10^{0}$ yr and $10^{6}$ yr. The two-level cloud loses energy to the BH and to infinity, primarily through quartic processes.
Since the 322 level grows through self-interactions to a very large occupation, its superradiant growth drives it to maximum occupation in just a few e-folds, thus reaching saturation much earlier than in the purely gravitational growth regime. As shown in the rightmost panels of Figure 2, for moderate self-interactions, 322 has grown maximally in $\sim 10^{6}$ yr (lower panels) compared to $\sim 10^{8}$ yr in the purely-gravitational growth regime (upper panels).
Once the quartic process $211\times 211\to 322\times{\rm BH}$ becomes faster than the superradiance of 211, the initial growth of 211 is suppressed by a simultaneous growth of 322. Quantitatively, the condition here comes about from Equation (9) by demanding $\gamma_{211\times 211}^{322\times{\rm BH}}(\varepsilon_{211}^{\max})^{2}\gtrsim 2\log(GM^{2})\gamma_{\rm BH}^{211}$. However, for small $\alpha$, the system equilibrates first, before this condition can be met. Setting $\dot{\varepsilon}_{211}=\dot{\varepsilon}_{322}=0$ in Equations (9) and (10) gives a different scaling for the threshold $f$ to transition to the moderate self-interaction regime for $\alpha\lesssim 0.04$. We refer the reader to Section IV.B.4 and Appendix E of [44] for the detailed computation. In both cases, we have a much smaller maximum occupation of the 211 level, $\varepsilon_{211}^{\max}\ll 1$, which occurs for
@@
Breakdown of perturbation theory: In all of the aforementioned regimes, the perturbative treatment of self-interactions is justified in the sense that the hydrogenic wavefunctions are not significantly distorted, timescales for all rates are much longer than oscillation times, and the dynamics are restricted to a two-level system. For $\alpha\gtrsim 0.22$, the two-level approximation no longer holds and higher levels can grow. In this case, additional signatures and transitions between higher levels can be present [44, 73]. While perturbative evolution can, in principle, be hierarchical and closed in non-gravitational regimes for $\alpha\lesssim 0.3$ [44], no demonstration of a closed regime which is under good analytic control has been performed for $\alpha\gtrsim 0.3$. We do not include these high-$\alpha$ scenarios in our current work as we only have control of our approximations in the purely gravitational growth regime, where $f$ is close to the Planck scale. There are potential signatures in the regime of $\alpha\gtrsim 0.22$, as shown in the results of this study, so we stress the importance of studying the high-$\alpha$ dynamics and parameter space in the future.
@@ 3.2 Occupation numbers and gravitational-wave strain
The emission frequencies for the 211 and 322 annihilations are broadly similar for a given $m_{b}$, whereas the transition signal is at a much lower frequency. For annihilation signals, the GW frequency corresponds to twice the energy of the annihilated particles, whereas for transition signals, it is the difference between the energy levels. The energy of the 211 and 322 levels differs only to the second order in $\alpha$ [see Equation (2.1)], and thus for small $\alpha$ as considered here, the energy of the 211 and 322 levels differ by a small relative amount. This leads to similar annihilation frequencies but a much lower transition energy difference than the total particle energy, leading to low-frequency transition GWs.
@@ 4.2 Observable parameter space
Figure 7: Parameter space reachable by current and next-generation detectors (enclosed by the solid curves; see legend) for (left) 211 annihilation, (middle) 322 annihilation, and (right) $322\rightarrow 211$ transition signals from an ultralight scalar cloud around Cygnus X-1, assuming parameters as in Table 1 and an age of $1.0\times 10^{5}$ yr. The dashed curves divide the parameter spaces into three regimes: (top) the strong self-interaction regime, (middle) the moderate self-interaction regime, and (bottom) the gravitational regime. The reachable regions for Einstein Telescope (enclosed by black curve) and Cosmic Explorer (enclosed by red curve) largely overlap. The parameter space reachable by MAGIS-Space is limited to the brightest spot in the right panel in this example. In other cases it can probe a larger parameter space; see for, e.g., MOA-2011 in Figure 11. For reference, the QCD axion’s mass to $1/f$ relation is shown by a dotted white line (in the bottom right corner of each plot). Axion interactions with Standard Model photons, electrons, and nuclei result in exclusions in typical models of GeV/$f\gtrsim 10^{-8}-10^{-9}$[104], complementary to the region shown here. See also [44, 45] for constraints from BH spin measurements.
@@ 5 Conclusions
We note that in this work we have only considered central measured values on some BH parameters and assumptions on others (like the initial spin and age) that have large uncertainties, so our projections on the reachable parameter space cannot be considered as true exclusions in the scalar particle parameter space. Data-driven constraints obtained from GW searches marginalized over the BH uncertainties would be of great interest. In particular, more accurate and precise measurements of the BH properties, e.g., mass, spin, age, and orbital parameters (if relevant), with future advanced telescopes will lead to smaller uncertainties in the estimates of the signal strength. By comparing the search sensitivities (with particular false alarm and false dismissal probabilities) to the estimated signals by taking into consideration the signal uncertainties propagated from the BH properties, one can obtain robust constraints on the existence of ultralight scalars.
@@ C.1 Cygnus X-1 with an age of $10^{6}$ yr
Figure 10: (Similar to Figure 7) Parameter space reachable by current and next-generation detectors (enclosed by the solid curves; see legend) for signals from Cygnus X-1, assuming parameters as in Table 1 and an age of $1.0\times 10^{6}$ yr. The reachable regions for the 211 annihilation, 322 annihilation, and $322\rightarrow 211$ transition signals are shown in the left, middle, and right panels, respectively. The dashed curves divide the parameter spaces into three regimes: (top) the strong self-interaction regime, (middle) the moderate self-interaction regime, and (bottom) the gravitational regime. The reachable regions for Einstein Telescope (enclosed by black curve) and Cosmic Explorer (enclosed by red curve) largely overlap. The transition signal is not reachable by MAGIS-Space for this particular example. For reference, the QCD axion’s mass to $1/f$ relation is shown by a dotted white line (in the bottom right corner of each plot).